Sabas Company has 20,000 shares of $100 par, 2% cumulative preferred stock and 100,000 shares of $50 par common stock. The following amounts were distributed as dividends:

Year 1 $10,000
Year 2 45,000
Year 3 90,000


Determine the dividends in arrears for preferred stock for the second year.

a.$10,000

b.$30,000

c.$0

d.$25,000
Answer #1

Arrears of preferred dividend is carried over next years and paid before paying any dividend on common shares

Annual preferred dividend = 20,000*100*2% = $40,000

Arrears in Year 1 = 40,000-10,000 = $30,000

Year 2 = 45000 - 40,000-30,000 = $25000

Hence, the answer is d.